The Standford -Binet IQ test is scaled so that he scores have a mean of 100...

  1. The Standford -Binet IQ test is scaled so that he scores have a mean of 100 and a standard deviation of 15.

    What is the percentage of scores that are less than 115?
Answer #1
for normal distribution z score =(X-μ)/σ
here mean=       μ= 100
std deviation   =σ= 15.0000

percentage of scores that are less than 115 :

probability = P(X<115) = P(Z<1)= 0.8413~ 84.13%